# Scalability of Metropolis-within-Gibbs schemes for high-dimensional Bayesian models

### v1

@@ D.7 Proof of Lemma 2
[…]
=$M\biggl{\lvert}\int_{\mathcal{X}}\int_{\mathcal{X}_{i}}g(\textbf{x},y_{i})\pi_{1,i}(\text{d}x_{i}\mid\textbf{x}_{-i})&\pi_{1,i}(\text{d}y_{i}\mid\textbf{x}_{-i})\pi_{1}(\text{d}\textbf{x}_{-i})\\ &-\int_{\mathcal{X}}\int_{\mathcal{X}_{i}}g(\textbf{x},y_{i})\pi_{1,i}(\text{d}x_{i}\mid\textbf{x}_{-i})\pi_{2,i}(\text{d}y_{i}\mid\textbf{x}_{-i})\pi_{1}(\text{d}\textbf{x}_{-i})\biggr{\rvert}\end{aligned}$
@@ D.9 Proof of Lemma 3
Denote with Gap$(P)$ the spectral gap of the operator $P$ (see e.g. [2] for the definition). It is well-known (see e.g. Lemma $2$ in [45]) that Gap$(P)=\min_{j\in\{1,\dots,J\}}\,\text{Gap}(P_{j})$. By Cheeger’s bounds (e.g. Lemma 5 in [2]) we have
@@ D.10 Proof of Theorem 4
Moreover, by the Bernstein-von Mises Theorem (e.g. Theorem $10.1$ in [68]), whose assumptions are met thanks to $(B1)-(B3)$, we can deduce
@@
D.11 Proof of Corollary 2
Proof.
Combining Theorems 5 and 4 we get
as $J\to\infty$, with $g(s)>0$ for every $s>0$. Let now
then by Lemma 1 and (58) we conclude
as $J\to\infty$. ∎
D.12 Proof of Theorem 5
@@
Let $P$ be a Markov kernel which is reversible with respect to $\pi$. For every $s,\epsilon\in(0,1/2)$ we have
@@
We need moreover two other preliminary Lemmas. These can be seen as the analogue of Lemma $4.1$ and Theorem $4.2$ in [3] to the setting of random-scan Gibbs sampler, and the proofs follow similar lines. Let $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}$ be a product space and $\pi\in\mathcal{P}(\mathcal{X})$, whose associated Gibbs sampler has operator $G$ as in (22). Let $T\,:\,\mathcal{X}_{1}\,\to\,\hat{\mathcal{X}}_{1}$ such that
@@
Consider model (20) and the Gibbs sampler defined in (22). Then, under assumptions $(B1)$-$(B6)$, for every $\psi^{*}\in\mathbb{R}^{D}$, $M\geq 1$ and $\epsilon>0$, there exists $T(\psi^{*},\epsilon,M)$ such that
@@
[…]
@@ D.13 Proof of Proposition 4
Without loss of generality we can assume $S^{*}$ to be compact, which by Weierstrass’ Theorem implies that assumption $(C)$ is satisfied. Moreover, assumptions $(B4)-(B6)$ are satisfied by Lemma $5.3$ in [3], so that the result follows by Corollary 2. ∎
D.14 Proof of Proposition 5
@@
[…]
@@
where $\pi_{J}^{(-2)}$ and $\tilde{\pi}^{(-2)}_{J}$ are the marginal distributions of $\psi$ under $\pi_{J}$ and $\tilde{\pi}_{J}$, respectively. Thus, by the Bernstein-von Mises Theorem (e.g. Theorem $10.1$ in [68]), whose requirements are met thanks to assumptions $(B1)$-$(B3)$, we have

### v2

@@ F.8 Proof of Lemma 2
$\biggl\lvert$ $\int_{\mathcal{X}}f(\textbf{y})\mu G_{1,i}(\text{d}\textbf{y})-\int_{\mathcal{X}}f(\textbf{y})\mu G_{2,i}(\text{d}\textbf{y})\biggr\rvert$
$=\begin{aligned} M\biggl\lvert\int_{\mathcal{X}}\int_{\mathcal{X}_{i}}g(\textbf{x},y_{i})\pi_{1,i}(\text{d}x_{i}\mid\textbf{x}_{-i})&\pi_{1,i}(\text{d}y_{i}\mid\textbf{x}_{-i})\pi_{1}(\text{d}\textbf{x}_{-i})\\ &-\int_{\mathcal{X}}\int_{\mathcal{X}_{i}}g(\textbf{x},y_{i})\pi_{1,i}(\text{d}x_{i}\mid\textbf{x}_{-i})\pi_{2,i}(\text{d}y_{i}\mid\textbf{x}_{-i})\pi_{1}(\text{d}\textbf{x}_{-i})\biggr\rvert\end{aligned}$
@@ F.10 Proof of Lemma 3
[…]
@@ F.11 Proof of Theorem 4
Moreover, by the Bernstein-von Mises Theorem (e.g. Theorem $10.1$ in [78]), whose assumptions are met thanks to (B1)(B3)(B1)-(B3), we can deduce
@@
F.12 Proof of Theorem 5
This proof requires multiple steps, which we summarize here:
  1. 1.
    We show that the $s$-conductance of an arbitrary operator $P$ can be lower bounded by the associated mixing times (Lemma 8).
  2. 2.
    We prove that the mixing times of the Gibbs sampler $G$ are uniformly bounded in probability (Lemma 10). In order to do this, we exploit a sufficiency structure which is available for hierarchical models (Lemma 9).
  3. 3.
    Finally, we combine points $1.$ and $2.$ to provide a lower bound on the $s$ conductance of $G$.
@@ F.12 Proof of Theorem 5
Let $P$ be a $\pi$-reversible Markov kernel. For every $s,\epsilon\in(0,1/2)$ we have
@@
We need moreover two other preliminary Lemmas. These can be seen as the analogue of Lemma $4.1$ and Theorem $4.2$ in [4] to the setting of random-scan Gibbs sampler, and the proofs follow similar lines. Let $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}$ be a product space and $\pi\in\mathcal{P}(\mathcal{X})$, whose associated Gibbs sampler has operator $G$ as in (22). Let $T\,:\,\mathcal{X}_{1}\,\to\,\hat{\mathcal{X}}_{1}$, with $\hat{\mathcal{X}}_{1}$ with $\hat{\mathcal{X}}_{1}$ being an arbitrary measurable space and $T$ such that
@@
For example, in the case of model (20), $\mathcal{L}(\psi\mid\bm{\theta},Y_{1:J})=\mathcal{L}\left(\psi\mid\bm{T}(\bm{\theta}),Y_{1:J}\right)$, with $\bm{T}$ the sufficient statistics defined in Section E.
@@
Consider model (20) and the Gibbs sampler defined in (22). Then, under assumptions (B1)(B1)-$(B6)$, for every $\psi^{*}\in\mathbb{R}^{D}$, $M\geq 1$ and $\epsilon>0$, there exists $T(\psi^{*},\epsilon,M)$ such that
@@
Let $\bm{T}$ be as in Section E, $\hat{\pi}_{J}=\mathcal{L}(\bm{T}(\bm{\theta}),\psi\mid Y_{1:J})$ under $(\bm{\theta},\psi)\sim\pi_{J}$ and $\hat{G}_{J}$ be the kernel of the Gibbs sampler targeting $\hat{\pi}_{J}$. Then, by Lemma 9 it holds $t_{mix}(G_{J},\epsilon,M)=t_{mix}(\hat{G}_{J},\epsilon,M)$. Let $\tilde{\psi}$ and $\tilde{\bm{T}}$ be the one-to-one transformations defined in $(17)$ and $(19)$ of [4]. Denoting $\tilde{\pi}_{J}=\mathcal{L}(\tilde{\bm{T}},\tilde{\psi}\mid Y_{1:J})$, by an analogous version of Lemma $2.1$ in [4] we have $t_{mix}(\hat{G}_{J},\epsilon,M)=t_{mix}(\tilde{G}_{J},\epsilon,M)$, where $\tilde{G}_{J}$ is the operator of the Gibbs sampler on $\tilde{\pi}_{J}$. By Lemma $C.18$ in [4] we have that
@@ F.13 Proof of Corollary 2
Combining Theorems 5 and 4 we get
as $J\to\infty$, with $g(s)>0$ for every $s>0$. Let now
then by Lemma 1 and (66) we conclude
as $J\to\infty$. ∎
F.14 Proof of Proposition 4
Proof.
Without loss of generality we can assume $S^{*}$ to be compact, which by Weierstrass’ Theorem implies that assumption $(C)$ is satisfied. Moreover, assumptions (B4)(B6)(B4)-(B6) are satisfied by Lemma $5.3$ in [4], so that the result follows by Corollary 2. ∎
F.15 Proof of Proposition 5
@@
By Lemma $C.45$ in [4] there exist a constant $R=R(c)\in[1,\infty)$ such that
@@
where $\pi_{J}^{(-2)}$ and $\tilde{\pi}^{(-2)}_{J}$ are the marginal distributions of $\psi$ under $\pi_{J}$ and $\tilde{\pi}_{J}$, respectively. Thus, by the Bernstein-von Mises Theorem (e.g. Theorem $10.1$ in [78]), whose requirements are met thanks to assumptions (B1)(B1)-$(B3)$, we have